Find the mean of the first five whole numbers.
Find the mean of the first five whole numbers.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The first five Whole numbers are 0, 1, 2, 3, and 4.
Mean = (Sum of first five whole numbers)/ (Total number of whole numbers)
Then,
Sum of five whole numbers = 0 + 1 + 2 + 3 +4 = 10
Total Number of whole numbers = 5 Mean = (10/5)
=2
∴Mean of first five whole numbers is 2

Thanks